OpenStudy (anonymous):

differentiate the following and simplify your answer into the specified form: g(x)=x(7sin3x+2cos3x) Answer: g(x)=(Cx+D)sin3x+(Ex+F)cos3x with constants. Find constants C,D,E,F

OpenStudy (anonymous):

you mean \[g'(x)=(Cx+D)\sin3x+(Ex+F)\cos3x \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me see! i will write soon!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

c=3 , d=7 , e=-6 , f=2

OpenStudy (anonymous):

sorry i make a mistake!!!

OpenStudy (anonymous):

c=-6,d=7, e=3, f=2

OpenStudy (anonymous):

\[g'(x)=(-6x+7)\sin3x + (3x+2)\cos3x\] \[g′(x)=(Cx+D)\sin3x+(Ex+F)\cos3x\] compare both equations and you will get the answer.